Some of you noticed that even though the database form was available, data from Feb-7 to present were not available in the archive. This was a hangup in the archiving script, due to a partially downloaded file at the point the server went down. The archive is now fully available, and all data to present is searchable.
The telescope was down from 03:02 to 03:41 UT on 2012-02-24 to address an issue with the focus. The focus algorithm was adjusted.
SII filters (6717, and 6730) are now in filter positions 15 and 16, replacing the UV_continuum and OH filters. The filters now available are:
This information can also be found at RCT Filters